3r - 5s = -33
5r + 3s = 47

Answer by checkley77(12844)

3r - 5s = -33 MULTIPLY THIS EQUATION BY 3
5r + 3s = 47 MULTIPLY THIS EQUATION BY 5 & ADD THEM.
9R-15S=-99
25R+15S=235
----------------------
34R=136
R=136/34
R=4 ANSWER.
3*4-5S=-33
12-5S=-33
-5S=-33-12
-5S=-45
S=-45/-5
S=9 ANSWER.
PROOF:
5*4+3*9=47
20+27=47
47=47